OpenStudy (anonymous):

PLEASE HELP!!!!!! 0.6x+0.8=1.4 What does x =

OpenStudy (aripotta):

subtract 0.8 from both sides

OpenStudy (anonymous):

k now what

OpenStudy (aripotta):

now divide 0.6 from both sides

OpenStudy (anonymous):

can you type all the steps please

jimthompson5910 (jim_thompson5910):

0.6x+0.8=1.4 0.6x+0.8-0.8=1.4-0.8 ... Subtract 0.8 from both sides. 0.6x+0=0.6 0.6x=0.6 0.6x/0.6=0.6/0.6 ... Divide both sides by 0.6. 1x = 1 x = 1

OpenStudy (anonymous):

thanks